OpenStudy (anonymous):

What's 75:35 simplified?

OpenStudy (jdoe0001):

\(\bf \cfrac{\cancel{75}}{\cancel{35}}\implies ?\)

OpenStudy (pinklion23):

divide by5 and there you go

OpenStudy (igreen):

Divide the numerator and denominator by 5.

OpenStudy (igreen):

75 / 5 = ? 35 / 5 = ?

OpenStudy (pinklion23):

75/5=? 35/5=?

OpenStudy (anonymous):

That would be 15:7, right?

OpenStudy (pinklion23):

15:7 yep

OpenStudy (anonymous):

Thx

OpenStudy (pinklion23):

np

OpenStudy (pinklion23):

thx for the medal

OpenStudy (anonymous):

YW :)